### 22 Dec 2023 v1

# Cross-border Exchange of CBDCs using Layer-2 Blockchain

Abstract
This paper proposes a novel multi-layer blockchain architecture for the cross-border trading of CBDCs. The permissioned layer-2, by relying on the public consensus of the underlying network, assures the security and integrity of the transactions and ensures interoperability with domestic CBDCs implementations. Multiple Layer-3s operate various Automated Market Makers (AMMs) and compete with each other for the lowest costs. To provide insights into the practical implications of the system, simulations of trading costs are conducted based on historical FX rates, with Project Mariana as a benchmark. The study shows that, even with liquidity fragmentation, a multi-layer and multi-AMM setup is more cost-efficient than a single AMM.
keywords:
Decentralized Finance Automated Market Maker CBDC Layer-2 Blockchain
1 Introduction
The Bitcoin network, launched in 2008, was first described in a whitepaper by the anonymous individual or group, known as Satoshi Nakamoto [1]. The fundamental promise of this first cryptocurrency was to establish a peer-to-peer payment system that operates independently of traditional financial intermediaries. In 2014, Vitalik Buterin introduced Ethereum [2], the next generation blockchain with a virtual machine capable of executing computer programs known as smart contracts. With Ethereum and the Ethereum Virtual Machine (EVM), Decentralized Finance (DeFi) was born. By utilizing smart contracts, DeFi offers sophisticated financial services, such as trading, lending, borrowing, derivatives and asset management [3, 4, 5, 6].
It quickly became apparent that this new blockchain-based financial system required non-volatile tokens, giving the advent to stablecoins—tokens with values pegged to fiat currencies. The success of stablecoins prompted central banks to explore similar solutions, leading to the development of Central Bank Digital Currencies (CBDCs) [7]. While CBDCs do not necessarily have to be implemented on a blockchain, the decentralized approach promises enhanced scalability and security [8, 9, 10, 11]. Various central banks across the globe are already engaged in CBDC initiatives[8, 9, 12]. The collective effort explores ways to integrate these CBDCs into a unified marketplace, potentially transcending borders, as seen in Project Mariana by the Bank for International Settlements (BIS) [8].
This paper introduces a novel architectural framework for cross-border CBDC trading that is based on Layer-2 (L2) blockchain scaling. L2s proved to be an efficient approach for addressing the limitations of Ethereum, especially scalability and privacy. [13, 14, 15]. The proposed system is the rollup on the public L1 blockchain that seamlessly integrates additional Layer-3 (L3) blockchains, each operating decentralized exchanges (DEXes), or other DeFi protocols, utilizing CBDCs. The system automatically selects L3 and DEXes with optimal costs. This system architecture is compared to the approach of Project Mariana in a series of simulations based on the historical Foreign Exchange (FX) rates.
2 Background
This section introduces the concept of Layer-2 blockchain, presents the functionalities of Project Mariana, and summarizes this work’s contributions to the research on cross-border CBDCs exchange.
Figure 1: Architecture of Rollup—Layer-2 blockchain that stores state in underlying Layer-1 blockchain
Figure 1: Architecture of Rollup—Layer-2 blockchain that stores state in underlying Layer-1 blockchain
Figure 2: Project Mariana Architecture: Cross-Border CBDC Exchange using Cryptoswap Invariant AMM
2.1 Layer-2 Blockchains
There are two approaches to tackle the challenges of blockchain scalability and related high gas prices: Layer-1 and Layer-2 blockchains. Layer-1 scaling involves creating entirely new blockchains, e.g., with unique consensus mechanisms or block sizes. These chains operate with their own validators and infrastructure, requiring decentralization to win users’ trust. Layer-2 scaling takes a different approach and uses complex computations off-chain (outside of Ethereum), aiming to reduce on-chain data congestion. These off-chain activities include rollups, sidechains, plasma, and state channels. Layer-2 blockchains rely on the security of underlying Layer-1 chains, such as Ethereum, for final settlement.
Rollups represent a form of Layer-2 scaling that does not custody any data by themselves. Rollups offload complex calculations from the Ethereum mainnet and store the results (along with other transactions) in Ethereum after compressing them. Fig. 2 illustrates rollup architecture with key components—sequencers and verifiers. Sequencers roll up transactions to the Layer-1 chain. By bundling transactions, rollups manage to save on gas fees. Verifiers are smart contracts that operate on Ethereum and verify the transactions stored by the sequencer. They ensure the correctness of the transactions.
2.2 Project Mariana
The specific focus of this work is on facilitating cross-border exchanges of CBDCs. One notable proposal for such a solution is Project Mariana (L1-Mariana) [8], initiated by BIS, the Swiss National Bank, the Bank of France, and the Monetary Authority of Singapore.
Within L1-Mariana, a DEX with an Automated Market Maker (AMM) is employed to exchange wholesale CBDCs versions of Swiss Franc (CHF), Euro (EUR), and Singaporean Dollar (SGD). The selected AMM is the Stableswap Invariant, first introduced by Curve v2—the second largest DEX in terms of trading volumes in permissionless DeFi on Ethereum [16]. The setup involves bridges that facilitate the transition of wholesale CBDCs from domestic central bank blockchains to the chain where the Cryptoswap Invariant pool operates, as depicted in Fig. 2.
Whereas the project proved the feasibility of AMM-DEX for cross-border CBDCs exchange, several questions arise—e.g., the choice of the blockchain (and its scalability and security), the AMM and the pool types (3-token pools vs 2-token pool)—which this research addresses.
3 Related Work and Contribution
There is an abundance of literature discussing the implementation of CBDCs in various countries [9, 10, 11, 17, 12], but little research on the exchange between CBDCs in AMMs, except for Project Mariana [8]. Lipton and Sepp[18] study the implicit costs (slippage cost) for G-10 CBDCs exchange at AMM with Concentrated Liquidity (Uniswap v3).
Our work is the first research that takes the holistic approach to analyze the total costs of CBDCs exchange at various AMMs, including Stableswap Invariant (Curve v2), applied in Project Mariana. This study found that a multi-AMM set-up for CBDCs exchange on L2 is more cost-efficient than a single AMM of Project Mariana, despite the fragmentation of liquidity among various AMMs. Such a multi-AMM system on private L2 is further presented, and its behavior is evaluated via simulation based on the historical FX rates, Mariana Project set up as a benchmark.
This work makes the following contributions.
[…]
4 System Architecture
This paper proposes a novel approach for facilitating cross-border CBDCs exchange through a rollup (a form of L2 blockchain) on the public Layer-1 network, further enhanced with a Layer-3 (L3), as illustrated in Fig. 3.
Figure 3: General Architecture of Cross-Border CBDC Exchange using L3 and L2s on Public Layer-1 Network
Blockchain and VM Type.  The system operates as a L2 blockchain on the public L1 network, benefiting from the inherent security and infrastructure of the L1, including consensus mechanisms and miners/validators. Bitcoin, Ethereum or other L1 networks can be the underlying L1. The system also incorporates Virtual Machines (VMs) to execute smart contracts, with the Ethereum Virtual Machine (EVM) being our chosen VM due to its compatibility with most DeFi protocols.
L2 and L3s Responsibilities.  L2 is the data integration layer and its operations are managed by central institutions like BIS, focusing on interoperability with other CBDC blockchains and facilitating communication between L2 and L3s. The system supports multiple independent L3 blockchains, each with the authority to determine the deployment of decentralized exchanges (DEX) and Automated Market Makers (AMM). Furthermore, the deployment of other DeFi protocols is feasible within this framework. AMM and exchange logic of L3s is moved to the external entities that compete with each other.
Transaction Privacy and Compliance.  The L2 and L3 blockchains are private (permissioned networks). All participants must undergo KYC [20] procedures before entering the system. This ensures the system can maintain full transaction privacy, e.g. by leveraging zero-knowledge proofs technology [21], while adhering to regulatory requirements.
Proof-of-storage.  We propose utilizing proof of storage protocols to enhance the security, instead of bridges. Unlike bridges that often involve trust assumptions and centralization, the system employs a proof-of-storage approach. CBDC tokens are directly minted on L2, and proof-of-storage protocols rely on corresponding token reserves within country-specific CBDC blockchains.
Rule-based router to L3s.  At the core of the system is a Rule-based router hosted on L2. This router employs rules to select the optimal AMM-DEX and L3 for traders at any given moment. Trading fees are dependent on gas costs and trade expenses, and the router optimizes these parameters for each trade.
Integration with DeFi on Ethereum.  To ensure compatibility with regulations, the system integrates with DeFi through the privacy pools on the Ethereum blockchain, proposed by Buterin et al.[22]. Only profiles and users validated through these pools are permitted to trade CBDCs, with no additional restrictions or limitations imposed.
Comparison to Mariana Project.  A key distinction is its flexibility regarding AMMs, allowing external operators to make decisions about the types of AMMs or central limit order books they wish to operate on L3. Operators of L3s retain control over actions and DeFi protocols within their respective layers. Another advantage of the L2-based system with L3 enhancements is cost optimization and resistance to network congestion. The L2 router automatically directs traders to the L3 AMM-DEX with the lowest trade and gas fees.
5 Metrics and AMMs
The major metric to consider for evaluation of the cross-border CBDCs exchange is the cost of the CBDCs swap. In general, the total cost of DeFi service consists of three components: [23, 24]
$TotalCost=GasFee+DeFiExplicitFee+DeFiImplicitFee$ (1)
Gas Fees are paid to the blockchain operators for executing the transaction, and DeFi Explicit Fee to the DeFi protocol for the provided service. DeFi Implicit Fee are indirect fees specific to the design of the DeFi protocol. For tokens swaps at AMM-DEX, swap fees is the explicit fee charged by the DEX, whereas slippage (price impact) is the implicit fee: [23, 24]
$TotalSwapCost=GasFee+SwapFee+SlippageCost$ (2)
SwapFee is typically the percentage of the transaction volume, and Liquidity Providers (LPs) supply capital to the AMM’s liquidity pools in exchange for participation in swap fees. Slippage Cost is the change in exchange price caused by the transaction and is specific to the AMM type, the liquidity pool size and transaction volume. The next section presents the major AMM types and corresponding slippage calculation methods.
5.1 Automated Market Makers
Constant Function Market Makers are the dominant type of AMMs. They are equipped with a reserve curve function $F:\mathbb{R}^{N}_{+}\rightarrow\mathbb{R}$ that maps $N$ token reserves $x_{i}$ to a fixed invariant (i.e., constant) $k$. The AMM-DEX in DeFi with the highest trading volumes [16] are Curve v2 [25] and Uniswap v3 [26]. The Project Mariana involves Curve v2 AMM and a 3-token pool. The formulation of AMMs follows [27].
Constant Product with Concentrated Liquidity Market Maker (CLMM) - Uniswap v3
$(x_{1}+\frac{L}{\sqrt{S_{l}}})(x_{2}+L\cdot\sqrt{S_{u}})=L^{2}$ (3)
Stableswap Invariant - Curve v1
$K\cdot D^{N-1}\cdot\sum\limits_{i=1}^{N}x_{i}+\prod\limits_{i=1}^{N}x_{i}=K\cdot D^{N}+(\frac{D}{N})^{N}$ (4)
where $K=A\cdot\prod_{i=1}x_{i}\cdot D^{-N}\cdot N^{N}$.
Cryptoswap Invariant - Curve v2  As in equation (4), but with K defined as
$K=A\cdot\underset{\text{ =: }K_{0}}{\underbrace{\frac{A\cdot\prod_{i=1}x_{i}}{D^{N}}\cdot N^{N}}}\cdot\frac{\gamma^{2}}{(\gamma+1-K_{0})^{2}}$ (5)
5.2 Slippage at AMM-DEXs
Further, we consider 2-token, or 3-token pools and denote the reserves of CBDCs in the pools by $x_{t}$ for CHF, $y_{t}$ for EUR, and $z_{t}$ for SGD at time t. The FX spot rate CHF-EUR is $S_{t}$ and CHF-SGD is $P_{t}$. The token reserves $x_{0}$, $y_{0}$, $z_{0}$ at the moment of the pool creation are set by:
$x_{0}=N_{0},y_{0}=S_{0}N_{0},z_{0}=P_{0}N_{0}$ (6)
For a buy transaction of $\Delta y$ EUR in exchange for $\Delta x$ CHF at time t and the slippage is:
$Slippage(\Delta x,\Delta y,S_{t})=\frac{\Delta y/\Delta x}{S_{t}}-1$ (7)
In CLMM, LPs can define the price range $[S_{l},S_{u}]$, for which liquidity is provided. When the exchange price $S_{t}$ moves outside the range $[S_{l},S_{u}]$, LP does not earn any fees. For simplicity, we consider the range $[\frac{S_{t}}{\alpha},S_{t}\alpha]$, where $\alpha>1$. Slippage at CLMM can be expressed as [27]:
$Slippage(t)=\frac{\Delta y}{N_{0}*\sqrt{S_{0}*S_{t}}}(1-\frac{1}{\sqrt{\alpha}})$ (8)
For $n>3$, the closed formula for slippage at Cryptoswap Invariant does not exist [28]. The numerical approximation of the Newton-Raphson method is used in two steps. First, given Eq. 6, the parameter D is calculated, and, subsequently, $\Delta x$ is found.
6 Simulation
To ensure comparability with L1-Mariana , we consider three CBDCs: Swiss Franc (CHF), Euro (EUR), and Singaporean Dollar (SGD) and back-test the total costs of CBDCs swaps in L1-Mariana and L2-L3-Exchange systems. We consider the buy transaction of EUR for CHF for volumes ranging from 1 EUR to 1mn EUR. The slippage costs for CLMMs are calculated based on Eq. 8, and the Newton-Raphson method is applied for Cryptoswap Invariants, both for $n=2$ and $n=3$. The implementation of all methods is available in a public GitHub repository [19] and was based on the smart contract code of Curve v2.
6.1 Data and Parameters
The analysis uses the historical exchange rates of CHF-EUR and CHF-SGD for the last three years (daily closing values from Yahoo Finance). We assume that arbitrageurs maintain prices aligned between CBDCs in AMMs and the FX market, which allows us to calculate the historical AMM pools’ compositions based on these FX rates.
Liquidity pools.  We assume 100mn CHF is provided to each system for liquidity provisions. As L1-Mariana operates Cryptoswap Invariant with one 3-token pool CHF-EUR-SGD, the pool’s value is 100mn CHF. In L2-L3-Exchange, we consider three L3 operators, each running AMM-DEX setup. The 100mn CHF liquidity is equally divided among the L3 operators. The first L3 operator uses a Cryptoswap Invariant with CHF-EUR-SGD pool (like in L1-Mariana), and the pool is valued 100/3 mn CHF. The second L3 also employs Cryptoswap Invariant, but with two 2-token pools: CHF-EUR and CHF-SGD, each valued 100/6 mn CHF. The third L3 operator utilizes CLMM with CHF-EUR and CHF-SGD pools, each valued 100/6 mn CHF. The L2-L3-Exchange pool architecture is illustrated in Fig. 7 in the appendix.
AMMs.  To be compatible with L1-Mariana, for all Cryptoswap Invariant we set A to 50 and gamma to $10^{-8}$. For CLMM, we consider the liquidity range around the spot price with the alpha of 1.2, rebalanced daily.
Swap fees.  Each liquidity pool charges swap fee of 1bp (i.e., 0.0001) of transaction volume.
Gas costs.  Gas fees for swaps on Ethereum, referred to in the Project Mariana, oscillated between 5 EUR and 50 EUR over the last three years [29]. The gas costs on L2s are, on average, 50 times lower compared to the underlying L1 [15, 14]. Based on these facts, we consider 15 EUR gas fees for a swap on L1 and 15EUR/50 on the L2-L3 network.
6.2 Results
Figure 4: Difference between the total swap costs at L1-Mariana and L2-L3-Exchange. The negative difference indicate L2-L3-Exchange out-performance for small and large transaction volumes.
Figure 5: Breakdown of total swap costs for various transaction volumes
Figure 6: Slippage costs for L1 Mariana (100mn TVL) and L3 AMM operators (each 33mn TVL)
The simulation outcomes are depicted in Figs. 4, 5 and 6. Fig. 8 in the appendix presents the historical total swap costs for transaction volumes of 10k, 100k, and 1mn EUR for CHF for each day over the last 3 years with fixed gas fees. Fig 4 extends this analysis for various gas fees and presents the difference between the costs of L2L3 Exchange and L1 Mariana. Notably, L2-L3-Exchange outperforms for volumes of 10k and 1mn. This out performance varies, influenced not only by transaction volume but also by current gas fees, as shown in fig. 4. For gas fees surpassing 800 EUR, L2-L3-Exchange outperforms for any transaction volume. For lower gas fees, such as 15 EUR, the out performance range is observed for volumes representing less than 0.01% or more than 0.7% of the provided liquidity. The reason for L2-L3-Exchange outperforming for smaller transactions lies in the minimization of gas fees through the application of L2 and L2s. Fig. 5 breaks down total swap costs, indicating that gas fees dominate the costs for L1 Mariana in smaller transactions, compared to the swap fee for L2-L3-Exchange. For larger transactions, slippage becomes the predominant component of total swap costs.
The out performance of the L2-L3-Exchange over L1-Mariana and its ability to handle slippage effectively stem from the use of L3 with CLMM, as illustrated in Figs. 6. The analysis of L3 selection indicates the use of L3 with Cryptoswap Invariant 2-token pool for smaller transactions, while the 3-token pool of Cryptoswap Invariant is unused for any transaction volumes. The simulation showed that the application of more AMMs, despite the liquidity fragmentation, outperforms the application of just one AMM: 100mn to L1-Mariana and 33.33mn to each L3. Fig. 9 simulates the swap costs when various total liquidity is provided.
7 Discussion
The primary consideration when establishing cross-border CBDCs exchange is the choice of blockchain architecture. While private L1s may initially appear to provide greater operational control, they require significant investments in physical infrastructure. Public blockchains operate in a decentralized manner, supported by well-established consensus mechanisms. A novel approach involves permissioned L2s networks on public blockchains, which combine the advantages of both private and public networks.
Another system decision is about allowance arbitrage, where the revenue from arbitrage could be allocated to AMM-DEX operators, L2-L3 operators or central banks. There is ongoing research about sequencer decentralization at public L2 blockchains and the enablement of MEV and MEV arbitrage. Currently, Ethereum roll-ups operations are supported only by one sequencer which does not allow MEV.
To assess the system, additional metrics such as latency and finality can be considered in conjunction with the domestic CBDC blockchains and the underlying L1 network. In this context, latency refers to the duration between submitting a transaction to a network and receiving the initial confirmation of its commitment. On the other hand, finality refers to the time required for confirming that a transaction is irrevocable.
Furthermore, the profitability of liquidity provisions and corresponding swap fee, should be analyzed. The prototype implementation of such systems could be based on the ready rollup frameworks [30, 31, 32].
8 Conclusions
The major finding of this research is that a multi-AMM setup on L2 for a CBDCs exchange is more cost-efficient than a single L1 AMM, despite the liquidity fragmentation. This work presented the design of such a multi-AMM system on private L2. The proposed system offers the advantages of private blockchain while minimizing infrastructure costs and enhancing security, as it leverages the consensus mechanism of the underlying L1.
Through a historical simulation based on the FX rates among CHF, EUR, and SGD, we compared the exchange costs between L2-L3 Exchange and Project Mariana. We found that the L2-L3 Exchange outperforms across small, low-medium, and large transactions, even in the presence of liquidity fragmentation between L3s. During gas fee spikes, the performance gap widens, with lower transaction costs for all transactions at L2-L3 Exchange. Our analysis of CBDC swap costs revealed that the gas fees component is significantly lower in the L2-L3 Exchange for medium transactions, enabling their efficient swap.
[…]
Acknowledgements
The first author is a research fellow at Matter Labs. The second author is a researcher at Matter Labs. This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.